# KinetiX: A performance portable code generator for chemical kinetics and transport properties

### v1

@@ A.2 Chemical kinetics
The third-body concentration for the $i$-th reaction $[X]_{i}$, reduced pressure $P_{r,i}$, and falloff blending factor $F_{i}$ are defined in the following sections.
@@
$\frac{\Delta S^{\circ}_{i}}{\mathcal{R}}$ $=$ $\sum_{k=1}^{N_{s}}\nu_{k,i}\frac{S^{\circ}_{k}}{\mathcal{R}}$ (16)
$\frac{\Delta H^{\circ}_{i}}{\mathcal{R}T}$ $=$ $\sum_{k=1}^{N_{s}}\nu_{k,i}\frac{H^{\circ}_{k}}{\mathcal{R}T}$ (17)
@@
$K_{p,i}=\exp\left(\sum_{k=1}^{N_{s}}\nu_{k,i}\left(\frac{S^{\circ}_{k}}{\mathcal{R}}-\frac{H^{\circ}_{k}}{\mathcal{R}T}\right)\right)=\exp\left(\sum_{k=1}^{N_{s}}\nu_{k,i}\frac{-G^{\circ}_{k}}{\mathcal{R}T}\right)$ (18)
@@ A.2.2 Falloff reactions
$k_{f,i}(T,P_{r,i})=k_{\infty,i}\left(\frac{P_{r,i}}{1+P_{r,i}}\right)F_{i}(T,P_{r,i})$ (24)
@@ A.3.1 Pure species viscosity
$\delta^{*}_{k}=\frac{1}{2}\frac{\mu_{k}^{2}}{\epsilon_{k}\sigma_{k}^{3}}.$ (34)
@@ A.3.2 Binary diffusion coefficients
$\epsilon_{np}$ $=$ $\xi^{2}\sqrt{\epsilon_{n}\epsilon_{p}}$ (42)
$\sigma_{np}$ $=$ $\frac{1}{2}(\sigma_{j}+\sigma_{k})\xi^{-\frac{1}{6}}$ (43)
$\mu^{2}_{np}$ $=$ 00 0 (44)

### v2

@@ A.2 Chemical kinetics
The third-body concentration for the $i$-th reaction $[X]_{i}$, reduced pressure $P_{r,i}$, and falloff blending factor $F_{i}$ is defined in the following sections.
@@
$\frac{\Delta S^{\circ}_{i}}{\mathcal{R}}$ $=$ $\sum_{k=1}^{N_{s}}\nu_{k,i}\frac{S^{\circ}_{k}}{\mathcal{R}},$ (16)
$\frac{\Delta H^{\circ}_{i}}{\mathcal{R}T}$ $=$ $\sum_{k=1}^{N_{s}}\nu_{k,i}\frac{H^{\circ}_{k}}{\mathcal{R}T},$ (17)
@@
$K_{p,i}=\exp\left(\sum_{k=1}^{N_{s}}\nu_{k,i}\left(\frac{S^{\circ}_{k}}{\mathcal{R}}-\frac{H^{\circ}_{k}}{\mathcal{R}T}\right)\right)=\exp\left(\sum_{k=1}^{N_{s}}\nu_{k,i}\frac{-G^{\circ}_{k}}{\mathcal{R}T}\right).$ (18)
@@ A.2.2 Falloff reactions
$k_{f,i}(T,P_{r,i})=k_{\infty,i}\left(\frac{P_{r,i}}{1+P_{r,i}}\right)F_{i}(T,P_{r,i}),$ (24)
@@
A.2.3 Pressure dependent reactions
For certain reactions, the pressure dependence cannot be adequately described using the modification factor $c_{i}$ (Eq. (8)) and falloff approach outlined previously. In these cases, an alternative formulation based on logarithmic interpolation between two reference pressures can be employed [46, 9]. At each reference pressure, the rate constant follows the modified Arrhenius expression:
$k_{1}(T)$ $=$ $A_{1}T^{\beta_{1}}\exp\left(-\frac{E_{\mathcal{R},1}}{T}\right)\quad\text{at}\quad p_{1}\quad\text{and}$ (32)
$k_{2}(T)$ $=$ $A_{2}T^{\beta_{2}}\exp\left(-\frac{E_{\mathcal{R},2}}{T}\right)\quad\text{at}\quad p_{2},$ (33)
where the Arrhenius parameters $(A_{1},\beta_{1},E_{\mathcal{R},1})$ and $(A_{2},\beta_{2},E_{\mathcal{R},2})$ are specified at pressures $p_{1}$ and $p_{2}$ respectively. For any intermediate pressure $p$ between $p_{1}$ and $p_{2}$, the forward rate constant can then be determined through logarithmic interpolation:
$\log k_{f}(T,p)=\log k_{1}(T)+(\log k_{2}(T)-\log k_{1}(T))\frac{\log p-\log p_{1}}{\log p_{2}-\log p_{1}}.$ (34)
@@ A.3.1 Pure species viscosity
$\delta^{*}_{k}=\frac{1}{2}\frac{\mu_{k}^{2}}{\epsilon_{k}\sigma_{k}^{3}}$ (37)
@@ A.3.2 Binary diffusion coefficients
$\epsilon_{np}$ $=$ $\xi^{2}\sqrt{\epsilon_{n}\epsilon_{p}},$ (45)
$\sigma_{np}$ $=$ $\frac{1}{2}(\sigma_{j}+\sigma_{k})\xi^{-\frac{1}{6}},$ (46)
$\mu^{2}_{np}$ $=$ $0,$ (47)